You want to withdraw $6,500 once year at the end of each year for the next 11years. r=8%. How much money do you need to deposit now to fund the account fully?
Answer #1

Amount to be deposited now=Amount withdrawn per year/rate*(1-1/(1+rate)^number of years)=6500/8%*(1-1/(1+8%)^11)=46403.26768